the two consecutive even integers whose product is 624
Answer by mananth(16946)

let the numbers be x,& x+2
x(x+2)=624
x^2+2x=624
x^2+2x-624=0
x^2+26x-24x-624
x(x+26)-24(x+26)=0
(x-24)(x+26)=0
x= 24 Or x = -26
the numbers are (24, 26) OR (-24-26)